Circular No. 6542

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) Z. Sekanina, Jet Propulsion Laboratory; and H. Böhnhardt, University

of Munich, report: “The comet’s persisting porcupine-like appearance (cf. IAUC 6463), consisting of 6–8 nearly-straight jets in various directions, is conceptually understood as a product of dust emission from 3–4 discrete active sources on the rotating nucleus. The jet pairs (one pair per source) are interpreted as boundaries of fan-shaped formations described by dust ejected from the sources continually between local sunrise and sunset (or continuously, if for any of the sources should the sun be above horizon throughout the day), a model that is consistent with evidence for more material being located in between the jets than on their outside, as seen on unprocessed images. One of several tentative solutions suggested by our computer simulation experiments that fit the observed jet-orientation pattern on several R-filter images taken with the large telescopes at the European Southern Observatory in Aug.–Nov. 1996, places the sunlit rotation pole near α = 30^{o}, δ = +45^{o} (equinox 2000.0), which implies an obliquity of the comet’s orbit plane to its equatorial plane of 140^{o}. The three most active areas have polar distances of ∼ 20^{o}, 55^{o}, and 130^{o}. These results are insensitive to both the comet’s rotation sense and its spin rate.”

M. Womack and D. Faith, Pennsylvania State University at Erie; M. C. Festou, Observatoire Midi-Pyrenees, Toulouse; and D. Slater and S. A. Stern, Southwest Research Institute, Boulder, report observations of comet C/1995 O1 with the NRAO 12-m telescope at Kitt Peak: “The J_{K}=3_{03}- 2_{02} transition of H_{2}CO at 218 GHz was detected on 1996 Dec. 15 and 1997 Jan. 17 UT with an average line flux area of 0.168 ± 0.008 K km/s. A production rate of Q(para-H_{2}CO) = 2.6 × 10^{26} molecules/s is derived, assuming a rotational temperature of 30 K and expansion velocity of 0.6 km/s. Comparison with observations of the J_{K}=3_{12}-2_{11} ortho transition of H_{2}CO at 226 GHz on 1996 Dec. 14, 15, and 1997 Jan. 17 indicate a weighted average ortho-to-para H_{2}CO ratio of 1.5 ± 0.3 for C/1995 O1 over the range of heliocentric distances 1.6–2.0 AU. This ratio is lower than that reported for H_{2}O in comets 1P/Halley and C/1986 P1 (Wilson), is similar to what is observed for H_{2}CO in cold, dense interstellar clouds, and may have implications for the thermal history of C/1995 O1.”

SUPERNOVAE Corrigendum. On IAUC 6540, line 6, for indicates the discovery tele-

scope: read indicates the telescope used to obtain the redshifts:

1997 January 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green